OpenStudy (anonymous):

solve for completing the square. x squared -8x+12=0. show work please :)

OpenStudy (anonymous):

*by

OpenStudy (lgbasallote):

x^2 - 8x +12 = 0 x^2 - 8x = -12 x^2 - 8x +(8/2)^2 = -12 + (8/2)^2 x^2 - 8x + (4)^2 = -12 + (4)^2 x^2 - 8x + 16 = -12 + 16 (x-4)^2 = 4 x-4 = sqrt (4) x - 4 = +/- 2 x = 4 +/- 2 x = 6 x = 2

OpenStudy (anonymous):

x squared -8x +12 = 0 -12 -12 x squared -8x + ____ = -12 + ____ (-8/2)squared -4squared 16 x squared -8x + 16 = -12 +16 (x-4)squared=4 (x-4)=+/-2 +4 +4 x=4+/- 2 x=6 x=2

OpenStudy (anonymous):

Ah! Beat me! ;)

OpenStudy (lgbasallote):

nice game :) you put up a good fight hehe

OpenStudy (ash2326):

@ineedhelppleaseee Whenever you need to complete square and you have say \[x^2+bx+c=0\] Just add and subtract \(\frac{b^2}{4}\) We get now \[x^2+bx+\frac{b^2}{4}-\frac{b^2}{4}+c=0\] We'll have now \[(x+\frac{b}{2})^2=\frac{b^2}{4}-c\] Now you can solve by taking square root both sides \[(x+\frac{b}{2})=\sqrt{\frac{b^2}{4}-c}\] so we get \[x=\sqrt{(\frac{b^2}{4}-c})-\frac{b}{2}\] Always remember this and you can solve any quadratic

OpenStudy (ash2326):

Sorry you'll get in the last step \[x=\pm\sqrt{(\frac{b^2}{4}-c)}-\frac{b}{2}\]

OpenStudy (anonymous):

Hey ash - how do you create the vertical fractions?

OpenStudy (ash2326):

See this @TheFigure http://openstudy.com/study#/updates/4f257bb9e4b0a2a9c266f130

OpenStudy (lgbasallote):

and ash just showed you how to derive the quadratic formula :DDD

OpenStudy (anonymous):

As long as a=1, yes. :)

OpenStudy (anonymous):

Thanks Ash!